A gas system has an initial temperature of 318.1k with the volume unknown. When the temperature changes to 53.10k, the volume is found to be 1270mL. What was the initial volume in mL


1
Expert's answer
2021-06-18T07:20:10-0400

V1/T1 = V2/T2

V1/318.1 = 1270/53.1

V1= 23.92 × 318.1

= 7608.04ml